# A Novel Multiscale Spatial Phase Estimate with the Structure Multivector

### v1

@@ Appendix 0.B Orientation Estimation with the SMV
$(M_{0}+M_{12}I_{2})^{2}=(e_{1}{\boldsymbol{n}})^{4}\left[A^{2}\cos^{2}({\boldsymbol{n}}\cdot{\boldsymbol{x}})+B^{2}\exp(4\epsilon I_{2})\cos^{2}({\boldsymbol{n}}^{\perp}_{\epsilon}\cdot{\boldsymbol{x}})+2AB\exp(2\epsilon I_{2})\cos({\boldsymbol{n}}\cdot{\boldsymbol{x}})\cos({\boldsymbol{n}}^{\perp}_{\epsilon}\cdot{\boldsymbol{x}})\right].$
@@
$\begin{split}(M_{1}+M_{2}I_{2})(M_{31}-M_{23}I_{2})&=(e_{1}{\boldsymbol{n}})^{4}[A^{2}\sin^{2}({\boldsymbol{n}}\cdot{\boldsymbol{x}})+B^{2}\exp(4\epsilon I_{2})\sin^{2}({\boldsymbol{n}}^{\perp}_{\epsilon}\cdot{\boldsymbol{x}})+\\ &AB(\exp(3\epsilon I_{2})+\exp(\epsilon I_{2}))\sin({\boldsymbol{n}}\cdot{\boldsymbol{x}})\sin({\boldsymbol{n}}^{\perp}_{\epsilon}\cdot{\boldsymbol{x}})].\end{split}$
Therefore our estimate is give by:
$\begin{split}\Theta(\epsilon)&=\theta+\frac{1}{4}\arg[A^{2}+B^{2}\exp(4\epsilon I_{2})+2AB\exp(2\epsilon I_{2})\cos({\boldsymbol{n}}\cdot{\boldsymbol{x}})\cos({\boldsymbol{n}}^{\perp}_{\epsilon}\cdot{\boldsymbol{x}})+\\ &AB(\exp(3\epsilon I_{2})+\exp(\epsilon I_{2}))\sin({\boldsymbol{n}}\cdot{\boldsymbol{x}})\sin({\boldsymbol{n}}^{\perp}_{\epsilon}\cdot{\boldsymbol{x}})]\\ &=\theta+\frac{1}{4}\tan^{-1}\left[\frac{B^{2}\sin(4\epsilon)+2AB\sin(2\epsilon)[\cos({\boldsymbol{n}}\cdot{\boldsymbol{x}})\cos({\boldsymbol{n}}^{\perp}_{\epsilon}\cdot{\boldsymbol{x}})+\cos(\epsilon)\sin({\boldsymbol{n}}\cdot{\boldsymbol{x}})\sin({\boldsymbol{n}}^{\perp}_{\epsilon}\cdot{\boldsymbol{x}})]}{A^{2}+B^{2}\cos(4\epsilon)+2AB\cos(2\epsilon)[\cos({\boldsymbol{n}}\cdot{\boldsymbol{x}})\cos({\boldsymbol{n}}^{\perp}_{\epsilon}\cdot{\boldsymbol{x}})+\cos(\epsilon)\sin({\boldsymbol{n}}\cdot{\boldsymbol{x}})\sin({\boldsymbol{n}}^{\perp}_{\epsilon}\cdot{\boldsymbol{x}})]}\right]\\ &\approx\theta+\frac{1}{4}\tan^{-1}\left[\frac{B^{2}\sin(4\epsilon)+2AB\sin(2\epsilon)\cos(({\boldsymbol{n}}-{\boldsymbol{n}}^{\perp}_{\epsilon})\cdot{\boldsymbol{x}})}{A^{2}+B^{2}\cos(4\epsilon)+2AB\cos(2\epsilon)\cos(({\boldsymbol{n}}-{\boldsymbol{n}}^{\perp}_{\epsilon})\cdot{\boldsymbol{x}})}\right].\end{split}$
We see that as $\frac{A}{B}\to 0$ it follows that $\Theta(\epsilon)\to\theta$.
Further, in the case $a=b$ we actually have

### v2

@@ Appendix 0.B Orientation Estimation with the SMV
$\begin{split}(M_{0}+M_{12}I_{2})^{2}&=(e_{1}{\boldsymbol{n}})^{4}\\ &[A^{2}\cos^{2}({\boldsymbol{n}}\cdot{\boldsymbol{x}})+B^{2}\exp(4\epsilon I_{2})\cos^{2}({\boldsymbol{n}}^{\perp}_{\epsilon}\cdot{\boldsymbol{x}})\\ &+2AB\exp(2\epsilon I_{2})\cos({\boldsymbol{n}}\cdot{\boldsymbol{x}})\cos({\boldsymbol{n}}^{\perp}_{\epsilon}\cdot{\boldsymbol{x}})].\end{split}$
@@
$\begin{split}(M_{1}+M_{2}I_{2})(M_{31}-M_{23}I_{2})&=(e_{1}{\boldsymbol{n}})^{4}[A^{2}\sin^{2}({\boldsymbol{n}}\cdot{\boldsymbol{x}})\\ &+B^{2}\exp(4\epsilon I_{2})\sin^{2}({\boldsymbol{n}}^{\perp}_{\epsilon}\cdot{\boldsymbol{x}})+AB(\exp(3\epsilon I_{2})\\ &+\exp(\epsilon I_{2}))\sin({\boldsymbol{n}}\cdot{\boldsymbol{x}})\sin({\boldsymbol{n}}^{\perp}_{\epsilon}\cdot{\boldsymbol{x}})].\end{split}$
In general for any real number $a$ we have: $\exp(ai)+\exp(3ai)=2\exp(2ai)\cos(a)$, our odd response becomes
$\begin{split}(M_{1}+M_{2}I_{2})(M_{31}-M_{23}I_{2})&=(e_{1}{\boldsymbol{n}})^{4}[A^{2}+B^{2}\exp(4\epsilon I_{2})\\ &+2AB\exp(2\epsilon I_{2})[\cos({\boldsymbol{n}}\cdot{\boldsymbol{x}})\cos({\boldsymbol{n}}^{\perp}_{\epsilon}\cdot{\boldsymbol{x}})+\\ &\cos(\epsilon)\sin({\boldsymbol{n}}\cdot{\boldsymbol{x}})\sin({\boldsymbol{n}}^{\perp}_{\epsilon}\cdot{\boldsymbol{x}})]].\end{split}$
Therefore, letting $M=\cos({\boldsymbol{n}}\cdot{\boldsymbol{x}})\cos({\boldsymbol{n}}^{\perp}_{\epsilon}\cdot{\boldsymbol{x}})+\cos(\epsilon)\sin({\boldsymbol{n}}\cdot{\boldsymbol{x}})\sin({\boldsymbol{n}}^{\perp}_{\epsilon}\cdot{\boldsymbol{x}})$ for brevity, our estimate is give by:
$\begin{split}\Theta(\epsilon)&=\theta+\frac{1}{4}\arg[A^{2}+B^{2}\exp(4\epsilon I_{2})+2AB\exp(2\epsilon I_{2})\cdot M]\\ &=\theta+\frac{1}{4}\tan^{-1}\left[\frac{B^{2}\sin(4\epsilon)+2AB\sin(2\epsilon)\cdot M}{A^{2}+B^{2}\cos(4\epsilon)+2AB\cos(2\epsilon)\cdot M}\right].\end{split}$
We see that as $\frac{B}{A}\to 0$ it follows that $\Theta(\epsilon)\to\theta$.
Additionally, in the case $A=B$, the argument in the artan function becomes
$\begin{split}\frac{\sin(4\epsilon)+2\sin(2\epsilon)\cdot M}{1+\cos(4\epsilon)+2\cos(2\epsilon)\cdot M}&=\frac{2\sin(2\epsilon)[\cos(2\epsilon)+M]}{2\cos(2\epsilon)[\cos(2\epsilon)+M]}\\ &=\tan(2\epsilon).\end{split}$
Thus